Events W,X,Y and Z are independent, and P(W)= 0.75, P(X)= 0.25, P(Y)= 0.2, and P(Z)= 0.3. Find P(W and X).
Answer by Edwin McCravy(20086)

Events W,X,Y and Z are independent, 
and P(W)= 0.75, P(X)= 0.25, P(Y)= 0.2,
and P(Z)= 0.3. Find P(W and X).

If they are independent, then "AND" means "MULTIPLY"

P(W and X) = P(W)·P(X) = (0.75)·(0.25) = 0.1875

Edwin